# Delzant type theorem for torus-equivariantly embedded toric hypersurfaces

### v1

@@ 2.3. Properties of the Rank of the Jacobian Matrices
[…]
@@ 3. Delzant Polytopes of Torus-equivariantly Embedded Toric Hypersurfaces
Hereafter, we assume that $j_{1}=1,\ldots,j_{n-1}=n-1$ otherwise specified.
@@ 3.1. Case of $\mathcal{J}_{\lambda}\neq\emptyset$
Let $\lambda$ be a vertex of the Delzant polytope $\Delta$. Define $\mathcal{J}_{\lambda}=\{i\mid\langle p_{1},v^{\lambda}_{i}\rangle=\cdots=\langle p_{n-1},v^{\lambda}_{i}\rangle=0\}$. In this section, we show that if the vertex $\lambda$ of $\Delta$ does not satisfy the condition (2) in Definition 1.1, then $\overline{C_{\lambda}(V)}$ is nonsingular.
@@
Let $j_{0}\in\mathcal{J}_{\lambda}$. By the definition of $\mathcal{J}_{\lambda}$, we obtain $v^{\lambda}_{j_{0}}\in V^{\perp}=\mathbb{R}q$. Since the vectors $v^{\lambda}_{j_{0}},q$ are nonzero, there exists $b\in\mathbb{R}\setminus\{0\}$ such that $q=bv^{\lambda}_{j_{0}}$. Since $\langle u^{\lambda}_{i},v^{\lambda}_{j}\rangle=\delta_{ij}$ (see [Yam24b, Lemma 3.10]), we have
@@
$-x=\sum_{j=1}^{n}s_{j}i_{V}^{\ast}(v^{\lambda}_{j})=\sum_{j=1}^{n-1}s_{j}i_{V}^{\ast}(v^{\lambda}_{j}).$
Let $R_{j}:=r_{j}+s_{j}$ for $j=1,\ldots,n$. Since $x+(-x)=0$, we obtain
$0=\sum_{j=1}^{n-1}R_{j}i_{V}^{\ast}(v^{\lambda}_{j}).$
Since we assume that $i_{V}^{\ast}(v^{\lambda}_{1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{n-1})$ are linearly independent, we obtain that $R_{1}=\cdots=R_{n-1}=0$. Since $r_{1},\ldots,r_{n-1},s_{1},\ldots,s_{n-1}\geq 0$, we obtain $r_{1}=\cdots=r_{n-1}=s_{1}=\cdots=s_{n-1}=0$, which means that $x=0$. Therefore, we obtain $W^{\lambda}_{V}=\{0\}$, i.e. there is no nontrivial subspace in the cone $i_{V}^{\ast}(\mathcal{C}_{\lambda})$. The point $i_{V}^{\ast}(\lambda)$ is a vertex of the convex polytope $i_{V}^{\ast}(\Delta)$. ∎
@@
Let $j\in\mathcal{J}_{\lambda}$, i.e. $i_{V}^{\ast}(v^{\lambda}_{j})=0$. From Lemma 3.1, the $(n-1)$ vectors $i_{V}^{\ast}(v^{\lambda}_{1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{j-1}),i_{V}^{\ast}(v^{\lambda}_{j+1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{n})$ are linearly independent.
@@ 3.2. Case of $i_{V}^{\ast}(\lambda)$ is not a vertex
Assume that the vectors $i_{V}^{\ast}(v^{\lambda}_{1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{n-1})$ are linearly independent. If $i_{V}^{\ast}(\lambda)$ is not a vertex of the polytope $i_{V}^{\ast}(\Delta)$, then $R_{n}\neq 0$ in Equation 3.1.
@@
Assume on the contrary that $R_{n}=0$. Then, from Equation 3.1 we obtain
@@
From Lemma 3.7, $R_{n}\neq 0$. Then, Equation 3.1 follows
@@
Assume on the contrary that $\langle u^{\lambda}_{n},q\rangle=0$. Then, from Equation 3.2, we obtain $\langle u^{\lambda}_{1},q\rangle=\cdots=\langle u^{\lambda}_{n-1},q\rangle=0$. Since $u^{\lambda}_{1},\ldots,u^{\lambda}_{n}$ form a $\mathbb{Z}$-basis of $\mathbb{Z}^{n}$, if $\langle u^{\lambda}_{1},q\rangle=\cdots=\langle u^{\lambda}_{n},q\rangle=0$, then $q=0$. This is contradiction to the assumption that $q$ is a basis of the orthogonal subspace $V^{\perp}$ to the $(n-1)$-dimensional subspace $V$ in $\mathfrak{t}^{n}\cong\mathbb{R}^{n}$. Therefore, $\langle u^{\lambda}_{n},q\rangle\neq 0$. ∎
@@
By Lemma 3.8, we obtain $\langle u^{\lambda}_{n},q\rangle\neq 0$. Moreover, since Equation 3.2 and Equation 3.3, if $n\in\mathcal{I}^{+}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}$, then $1,\ldots,n-1\in\mathcal{I}^{+}_{\lambda}$, i.e. $\mathcal{I}^{-}_{\lambda}\setminus\mathcal{I}^{0}=\emptyset$. Similarly, if $n\in\mathcal{I}^{-}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}$, then $1,\ldots,n-1\in\mathcal{I}^{-}_{\lambda}$, i.e. $\mathcal{I}^{+}_{\lambda}\setminus\mathcal{I}^{0}=\emptyset$. ∎
@@ 3.3. Case of $\mathcal{J}_{\lambda}=\emptyset$ and $i_{V}^{\ast}(\lambda)$ is a vertex
Assumption 3.11 implies that $i_{V}^{\ast}(\lambda)$ is a vertex of the polytope $i_{V}^{\ast}(\Delta)$. Note that if the polytope $i_{V}^{\ast}(\Delta)$ is simple, then Assumption 3.11 holds.
@@
00 0 $=\sum_{j=1}^{n}\langle u^{\lambda}_{j},q\rangle i_{V}^{\ast}(v^{\lambda}_{j})$
@@
Note that if $S_{j}=0$, then $\langle u^{\lambda}_{j},q\rangle=0$.
@@
from Equation 3.5. Since the vectors $u^{\lambda}_{1},\ldots,u^{\lambda}_{n}$ form a $\mathbb{Z}$-basis of $\mathbb{Z}^{n}$, Equation 3.6 means that $q=0$, i.e. $V^{\perp}=\{0\}$. This is contradiction to the assumption that $V^{\perp}$ is the orthogonal subspace to the $(n-1)$-dimensional subspace $V$ in $\mathfrak{t}^{n}\cong\mathbb{R}^{n}$. Therefore, $\langle u^{\lambda}_{n},q\rangle\neq 0$.
Since $\langle u^{\lambda}_{n},q\rangle\neq 0$, we obtain $n\not\in\mathcal{I}^{0}_{\lambda}$, i.e. $n\in\mathcal{I}^{+}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}$ or $n\in\mathcal{I}^{-}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}$. From Equation 3.5, if $n\in\mathcal{I}^{+}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}$, then $1,\ldots,n-1\in\mathcal{I}^{-}_{\lambda}$. Similarly, if $n\in\mathcal{I}^{-}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}$, then $1,\ldots,n-1\in\mathcal{I}^{+}_{\lambda}$. ∎
@@
Moreover, since $i_{V}^{\ast}(v^{\lambda}_{n})\in i_{V}^{\ast}(\mathcal{C}_{\lambda})$ from Assumption 3.11, $i_{V}^{\ast}(v^{\lambda}_{n})$ can be expressed as the form in Equation 3.4. From these equations, we obtain
@@
From Equation 3.5 and Equation 3.7, we obtain
@@
From Equation 3.8, we obtain
[${{}^{t}\!\left[\begin{matrix}u^{\lambda}_{1}&\cdots&u^{\lambda}_{n}\end{matrix}\right]}q=\left[\begin{matrix}\langle u^{\lambda}_{1},q\rangle\\ \vdots\\ \langle u^{\lambda}_{n},q\rangle\end{matrix}\right]=\langle u^{\lambda}_{n},q\rangle\left[\begin{matrix}-Q_{1}\\ \vdots\\ -Q_{n-1}\\ 1\end{matrix}\right]$
and since [${{}^{t}[u^{\lambda}_{1}\;\cdots\;u^{\lambda}_{n}]}[v^{\lambda}_{1}\;\cdots\;v^{\lambda}_{n}]=E_{n}$, we obtain
@@ 3.4. The Proof of the first Main Theorem
If the Delzant polytope $\Delta$ is good with respect to the map $i_{V}^{\ast}$, then the rank of the Jacobian matrix $Df^{\lambda}$ is equal to one at any point for any vertex $\lambda$ of $\Delta$. In particular, $\overline{C(V)}$ is a smooth complex hypersurface in the toric manifold $X$ corresponding to $\Delta$.
@@
Since we assume that the Delzant polytope $\Delta$ is good with respect to the map $i_{V}^{\ast}$, if a vertex $\lambda$ of $\Delta$ is a good vertex with respect to the map $i_{V}^{\ast}$, then Assumption 3.11 holds for such $\lambda$ and there exist distinct $j_{1},\ldots,j_{n-1}\in\{1,\ldots,n\}$ such that the vectors $i_{V}^{\ast}(v^{\lambda}_{j_{1}}),\ldots,i_{V}^{\ast}(v^{\lambda}_{j_{n-1}})\in\mathbb{Z}^{n-1}$ form a $\mathbb{Z}$-basis of $\mathbb{Z}^{n-1}$. In Corollary 3.14, we show that $\overline{C_{\lambda}(V)}$ is nonsingular in this case.
Therefore, if $\Delta$ is good with respect to the map $i_{V}^{\ast}$, then $\overline{C_{\lambda}(V)}$ is nonsingular for any vertex $\lambda$ of $\Delta$, i.e. $\overline{C(V)}$ is a smooth complex hypersurface in the toric manifold $X$. ∎
@@ 4. From Torus-equivariantly Embedded Toric Hypersurfaces to Delzant Polytopes
Assume that the rank of the Jacobian matrix $Df^{\lambda}$ is equal to one at any point of the zero locus of $f^{\lambda}$ for any vertex $\lambda$ of the Delzant polytope, i.e. $\overline{C(V)}$ is nonsingular. In this case, as we show in Proposition 2.5, $\overline{C(V)}$ is toric with respect to the Hamiltonian $T^{n-1}$-action defined in [Yam24b] with the moment map $\overline{\mu}=i_{V}^{\ast}\circ\mu\mid_{\overline{C(V)}}$. Moreover, we have the following:
Theorem 4.1 ([Yam24b, Theorem 4.20]).
@@
at the origin $o\in\varphi_{\lambda}(U_{\lambda})\cap\overline{C_{\lambda}(V)}$. If such $i_{0}\in\mathcal{I}^{+}_{\lambda}$, then we obtain $\langle u^{\lambda}_{j},q\rangle-\delta_{i_{0}j}=0$ for $j=1,\ldots,n-1$, i.e. $\langle u^{\lambda}_{i_{0}},q\rangle=1$ and $\langle u^{\lambda}_{j},q\rangle=0$ for any $j\in\mathcal{I}^{+}_{\lambda}\setminus\{i_{0}\}$. If such $i_{0}\notin\mathcal{I}^{+}_{\lambda}$, i.e. $i_{0}=n$, then $\langle u^{\lambda}_{n},q\rangle=-1$. ∎
@@
From Lemma 2.6, the left hand side of this equation is equal to zero. Therefore, we obtain Equation 4.1. ∎
@@
If $\lambda$ is good, Assumption 3.11 holds, and $\overline{C_{\lambda}(V)}$ is nonsingular, then we have $i_{V}^{\ast}(v^{\lambda}_{1})=-\langle u^{\lambda}_{n},q\rangle i_{V}^{\ast}(v^{\lambda}_{n})$ with $-\langle u^{\lambda}_{n},q\rangle\in\mathbb{Z}_{>0}$. Therefore, we may use the vector $i_{V}^{\ast}(v^{\lambda}_{n})$ instead of $i_{V}^{\ast}(v^{\lambda}_{1})$ in Assumption 3.11, i.e. we may assume the following:
@@
=±$=\pm\left(\langle q,v^{\lambda}_{1}\rangle+\langle q,v^{\lambda}_{n}\rangle\langle u^{\lambda}_{n},q\rangle\right)\det[i_{V}^{\ast}(v^{\lambda}_{2})\;\cdots\;i_{V}^{\ast}(v^{\lambda}_{n})].$
@@
±$\pm\left(\langle q,v^{\lambda}_{1}\rangle+\langle q,v^{\lambda}_{n}\rangle\langle u^{\lambda}_{n},q\rangle\right)\det[i_{V}^{\ast}(v^{\lambda}_{2})\;\cdots\;i_{V}^{\ast}(v^{\lambda}_{n})]$
=±$=\pm\langle q,q\rangle\det[i_{V}^{\ast}(v^{\lambda}_{2})\;\cdots\;i_{V}^{\ast}(v^{\lambda}_{n})].$
@@
Thus, we obtain Equation 4.2. ∎
@@
4.3. The Prood of the second Main Theorem
By calculating the determinant of the Gram matrix of $[p_{1}\;\cdots\;p_{n-1}\;q]$, we obtain the following:

### v2

@@ 2.3. Properties of the Rank of the Jacobian Matrices
$\frac{\partial f^{\lambda}}{\partial z^{\lambda}_{i}}=\begin{cases*}\langle u^{\lambda}_{i},q\rangle\prod_{j\in\mathcal{I}^{-}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}}(z^{\lambda}_{j})^{-\langle u^{\lambda}_{j},q\rangle-\delta_{ij}}&if $i\in\mathcal{I}^{-}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}$,\\ 0&otherwise,\end{cases*}$
@@ 3. Delzant Polytopes of Torus-equivariantly Embedded Toric Hypersurfaces
Hereafter, we assume that $j_{1}=1,\ldots,j_{n-1}=n-1$ otherwise specified. Let $\lambda$ be a vertex of the Delzant polytope $\Delta$. Define $\mathcal{J}_{\lambda}=\{i\mid\langle p_{1},v^{\lambda}_{i}\rangle=\cdots=\langle p_{n-1},v^{\lambda}_{i}\rangle=0\}$.
@@ 3.1. Case of $\mathcal{J}_{\lambda}\neq\emptyset$
In this section, we show that if the vertex $\lambda$ of $\Delta$ does not satisfy the condition (2) in Definition 1.1, then $\overline{C_{\lambda}(V)}$ is nonsingular.
@@
Let $j_{0}\in\mathcal{J}_{\lambda}$. By the definition of $\mathcal{J}_{\lambda}$, we obtain $v^{\lambda}_{j_{0}}\in V^{\perp}=\mathbb{R}q$. Since the vectors $v^{\lambda}_{j_{0}},q$ are nonzero, there exists $b\in\mathbb{R}\setminus\{0\}$ such that $q=bv^{\lambda}_{j_{0}}$. Since $\langle u^{\lambda}_{i},v^{\lambda}_{j}\rangle=\delta_{ij}$ (see [Yam25, Lemma 3.10]), we have
@@
$x=\sum_{j=1}^{n}s_{j}i_{V}^{\ast}(v^{\lambda}_{j})=\sum_{j=1}^{n-1}s_{j}i_{V}^{\ast}(v^{\lambda}_{j})=\sum_{j=1}^{n-1}(-r_{j})i_{V}^{\ast}(v^{\lambda}_{j}).$
Since we assume that $i_{V}^{\ast}(v^{\lambda}_{1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{n-1})$ are linearly independent, we obtain $s_{j}=-r_{j}$ for $j=1,\ldots,n-1$. Since $r_{1},\ldots,r_{n-1},s_{1},\ldots,s_{n-1}\geq 0$, we obtain $r_{1}=\cdots=r_{n-1}=s_{1}=\cdots=s_{n-1}=0$, which means that $x=0$. Therefore, we obtain $W^{\lambda}_{V}=\{0\}$, i.e. there is no nontrivial subspace in the cone $i_{V}^{\ast}(\mathcal{C}_{\lambda})$. The point $i_{V}^{\ast}(\lambda)$ is a vertex of the convex polytope $i_{V}^{\ast}(\Delta)$. ∎
@@
Let $j\in\mathcal{J}_{\lambda}$, i.e. $i_{V}^{\ast}(v^{\lambda}_{j})=0$. From Lemma 3.1, the (n1)(n-1) vectors $i_{V}^{\ast}(v^{\lambda}_{1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{j-1}),i_{V}^{\ast}(v^{\lambda}_{j+1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{n})$ are linearly independent.
@@ 3.2. Case of $i_{V}^{\ast}(\lambda)$ is not a vertex
Assume that the vectors $i_{V}^{\ast}(v^{\lambda}_{1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{n-1})$ are linearly independent. If $i_{V}^{\ast}(\lambda)$ is not a vertex of the polytope $i_{V}^{\ast}(\Delta)$, then $R_{n}\neq 0$ in Equation (3.1).
@@
Assume on the contrary that $R_{n}=0$. Then, from Equation (3.1) we obtain
@@
From Lemma 3.7, $R_{n}\neq 0$. Then, Equation (3.1) follows
@@
Assume on the contrary that $\langle u^{\lambda}_{n},q\rangle=0$. Then, from Equation (3.2), we obtain $\langle u^{\lambda}_{1},q\rangle=\cdots=\langle u^{\lambda}_{n-1},q\rangle=0$. Since $u^{\lambda}_{1},\ldots,u^{\lambda}_{n}$ form a $\mathbb{Z}$-basis of $\mathbb{Z}^{n}$, if $\langle u^{\lambda}_{1},q\rangle=\cdots=\langle u^{\lambda}_{n},q\rangle=0$, then $q=0$. This is contradiction to the assumption that $q$ is a basis of the orthogonal subspace $V^{\perp}$ to the (n1)(n-1)-dimensional subspace $V$ in $\mathfrak{t}^{n}\cong\mathbb{R}^{n}$. Therefore, $\langle u^{\lambda}_{n},q\rangle\neq 0$. ∎
@@
By Lemma 3.8, we obtain $\langle u^{\lambda}_{n},q\rangle\neq 0$. Moreover, since Equation (3.2) and Equation (3.3), if $n\in\mathcal{I}^{+}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}$, then $1,\ldots,n-1\in\mathcal{I}^{+}_{\lambda}$, i.e. $\mathcal{I}^{-}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}=\emptyset$. Similarly, if $n\in\mathcal{I}^{-}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}$, then $1,\ldots,n-1\in\mathcal{I}^{-}_{\lambda}$, i.e. $\mathcal{I}^{+}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}=\emptyset$. ∎
@@ 3.3. Case of $\mathcal{J}_{\lambda}=\emptyset$ and $i_{V}^{\ast}(\lambda)$ is a vertex
This assumption is the same as the first condition in Definition 1.2. Note that if the polytope $i_{V}^{\ast}(\Delta)$ is simple, then Assumption 3.11 holds.
@@
$0$ $=\sum_{j=1}^{n}\langle u^{\lambda}_{j},q\rangle i_{V}^{\ast}(v^{\lambda}_{j})$
@@
from Equation (3.5). Since the vectors $u^{\lambda}_{1},\ldots,u^{\lambda}_{n}$ form a $\mathbb{Z}$-basis of $\mathbb{Z}^{n}$, Equation (3.6) means that $q=0$, i.e. $V^{\perp}=\{0\}$. This is contradiction to the assumption that $V^{\perp}$ is the orthogonal subspace to the (n1)(n-1)-dimensional subspace $V$ in $\mathfrak{t}^{n}\cong\mathbb{R}^{n}$. Therefore, $\langle u^{\lambda}_{n},q\rangle\neq 0$.
Since $\langle u^{\lambda}_{n},q\rangle\neq 0$, we obtain $n\not\in\mathcal{I}^{0}_{\lambda}$, i.e. $n\in\mathcal{I}^{+}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}$ or $n\in\mathcal{I}^{-}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}$. From Equation (3.5), if $n\in\mathcal{I}^{+}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}$, then $1,\ldots,n-1\in\mathcal{I}^{-}_{\lambda}$. Similarly, if $n\in\mathcal{I}^{-}_{\lambda}\setminus\mathcal{I}^{0}_{\lambda}$, then $1,\ldots,n-1\in\mathcal{I}^{+}_{\lambda}$. ∎
@@
Moreover, since $i_{V}^{\ast}(v^{\lambda}_{n})\in i_{V}^{\ast}(\mathcal{C}_{\lambda})$ from Assumption 3.11, $i_{V}^{\ast}(v^{\lambda}_{n})$ can be expressed as the form in Equation (3.4). From these equations, we obtain
@@
From Equation (3.5) and Equation (3.7), we obtain
@@
From Equation (3.8), we obtain
${{}^{t}\!\left[\begin{matrix}u^{\lambda}_{1}&\cdots&u^{\lambda}_{n}\end{matrix}\right]}q=\left[\begin{matrix}\langle u^{\lambda}_{1},q\rangle\\ \vdots\\ \langle u^{\lambda}_{n},q\rangle\end{matrix}\right]=\langle u^{\lambda}_{n},q\rangle\left[\begin{matrix}-Q_{1}\\ \vdots\\ -Q_{n-1}\\ 1\end{matrix}\right]$
and since ${{}^{t}[u^{\lambda}_{1}\;\cdots\;u^{\lambda}_{n}]}[v^{\lambda}_{1}\;\cdots\;v^{\lambda}_{n}]=E_{n}$, we obtain
@@ 3.4. The Proof of the first Main Theorem
If $V$ is admissible with respect to $\Delta$, then the rank of the Jacobian matrix $Df^{\lambda}$ is equal to one at any point for any vertex $\lambda$ of $\Delta$. In particular, $\overline{C(V)}$ is a smooth complex hypersurface in the toric manifold $X$ corresponding to $\Delta$.
@@
Since we assume that $V$ is admissible with respect to $\Delta$, if a vertex $\lambda$ of $\Delta$ is a good vertex with respect to the map $i_{V}^{\ast}$, then Assumption 3.11 holds for such $\lambda$ and there exist distinct $j_{1},\ldots,j_{n-1}\in\{1,\ldots,n\}$ such that the vectors $i_{V}^{\ast}(v^{\lambda}_{j_{1}}),\ldots,i_{V}^{\ast}(v^{\lambda}_{j_{n-1}})\in\mathbb{Z}^{n-1}$ form a $\mathbb{Z}$-basis of $\mathbb{Z}^{n-1}$. In Corollary 3.14, we show that $\overline{C_{\lambda}(V)}$ is nonsingular in this case.
Therefore, if $V$ is admissible with respect to $\Delta$, then $\overline{C_{\lambda}(V)}$ is nonsingular for any vertex $\lambda$ of $\Delta$, i.e. $\overline{C(V)}$ is a smooth complex hypersurface in the toric manifold $X$. ∎
@@ 4. From Torus-equivariantly Embedded Toric Hypersurfaces to Delzant Polytopes
Assume that the rank of the Jacobian matrix $Df^{\lambda}$ is equal to one at any point of the zero locus of $f^{\lambda}$ for any vertex $\lambda$ of the Delzant polytope, i.e. $\overline{C(V)}$ is nonsingular. In this case, as we show in Proposition 2.5, $\overline{C(V)}$ is toric with respect to the Hamiltonian $T^{n-1}$-action defined in [Yam25] with the moment map $\overline{\mu}=i_{V}^{\ast}\circ\mu\mid_{\overline{C(V)}}$. Moreover, we have the following:
Theorem 4.1 ([Yam25, Theorem 4.20]).
@@
at the origin $o\in\varphi_{\lambda}(U_{\lambda})\cap\overline{C_{\lambda}(V)}$. If such $i_{0}$ is an element of $\mathcal{I}^{+}_{\lambda}$, then we obtain $\langle u^{\lambda}_{j},q\rangle-\delta_{i_{0}j}=0$ for $j=1,\ldots,n-1$, i.e. $\langle u^{\lambda}_{i_{0}},q\rangle=1$ and $\langle u^{\lambda}_{j},q\rangle=0$ for any $j\in\mathcal{I}^{+}_{\lambda}\setminus\{i_{0}\}$. If such $i_{0}\notin\mathcal{I}^{+}_{\lambda}$, i.e. $i_{0}=n$, then $\langle u^{\lambda}_{n},q\rangle=-1$. ∎
@@
From Lemma 2.6, the left-hand side of this equation is equal to zero. Therefore, we obtain Equation (4.1). ∎
@@
If $\lambda$ is good, Assumption 3.11 holds, and $\overline{C_{\lambda}(V)}$ is nonsingular, then we have $i_{V}^{\ast}(v^{\lambda}_{1})=-\langle u^{\lambda}_{n},q\rangle i_{V}^{\ast}(v^{\lambda}_{n})$ with $\langle u^{\lambda}_{n},q\rangle\in\mathbb{Z}_{>0}$. Therefore, we may use the vector $i_{V}^{\ast}(v^{\lambda}_{n})$ instead of $i_{V}^{\ast}(v^{\lambda}_{1})$ in Assumption 3.11, i.e. we may assume the following:
@@
$=\pm\left(\langle q,v^{\lambda}_{1}\rangle+\langle q,v^{\lambda}_{n}\rangle\langle u^{\lambda}_{n},q\rangle\right)\det[i_{V}^{\ast}(v^{\lambda}_{2})\;\cdots\;i_{V}^{\ast}(v^{\lambda}_{n})].$
@@
$\pm\left(\langle q,v^{\lambda}_{1}\rangle+\langle q,v^{\lambda}_{n}\rangle\langle u^{\lambda}_{n},q\rangle\right)\det[i_{V}^{\ast}(v^{\lambda}_{2})\;\cdots\;i_{V}^{\ast}(v^{\lambda}_{n})]$
$=\pm\langle q,q\rangle\det[i_{V}^{\ast}(v^{\lambda}_{2})\;\cdots\;i_{V}^{\ast}(v^{\lambda}_{n})].$
@@
Thus, we obtain Equation (4.2). ∎
@@
4.3. The Proof of the second Main Theorem
By calculating the determinant of the Gram matrix of $p_{1}\;\cdots\;p_{n-1}\;q]$, we obtain the following: